A child goes down a playground slide with an acceleration of 1.22 m/s2 . You may want to review (Pages 156 - 163) .

Part A

Find the coefficient of kinetic friction between the child and the slide if the slide is inclined at an angle of 34.0 ∘ below the horizontal.
Answer #1

given

a = 1.22 m/sec2

= 34.0

we have equation

fk = k N

= k ( m g cos )

and along the line the equation of motion is

m g sin - fk = m a

m g sin - k m g cos = m a

g sin - k g cos = a

k = ( g sin - a ) / g cos

= ( 9.8 x sin34 - 1.22 ) / 9.8 x cos34

= ( 5.48 - 1.22 ) / 8.124

= 4.26 / 8.124

k = 0.5243

so the coefficient of kinetic friction between the child and the slide is k = 0.5243